# Physics of droplet regulation in biological cells

### v3

@@ 5.5 More complex chemically active droplets
The behaviour of chemically active phase-separating systems can be much more complex than that described above when more species are involved. Since this topic represents a relatively new field of research, we list only a few recent discoveries here.
@@ 5.5.2 Self-organized drift
The combination of chemical turnover and diffusion generically creates concentration gradients, which can move particles and droplets by diffusiophoresis [483, 454, 189, 190, 484]. Concentration gradients also develop around chemically active droplets (see figure 5.3), implying that they can influence the positioning of other droplets. For instance, externally maintained droplets naturally repel each other, which stabilises their hexagonal stationary pattern. Their repulsion follows directly from the electrostatic analogy discussed in section 5.3.2.
@@ 5.5.3 Active multi-component mixtures
Active multi-component mixtures can exhibit various instabilities and patterns, ranging from simple Turing patterns [488, 445, 446, 472] to systems with multiple length scales [489], to more complex spatiotemporal behaviour [490, 491, 492]. Particularly interesting is the case where phase separation interferes with chemical reactions [493, 494], so that behaviours beyond those described above can emerge. For instance, droplets may now be stable even in the presence of passive reactions, which is impossible in binary systems [442]. In this case, internally maintained droplets do not fill the entire system, although their size typically still scales with system size [451, 435]. Externally maintained droplets, however, can either scale with system size or exhibit the size control discussed in section 5.3 [458, 495]. In both of these cases, the droplet size distribution relaxes to a monodisperse (and universal) distribution at late times [495]. Externally maintained droplets can also assume shell-like structures [453, 496], essentially because chemical reactions deplete the inside of the droplet so that concentration (see figure 5.3a) drops below the spinodal concentration and a bubble emerges. Chemically active droplets may be further controlled by enzyme positioning, either because they are mobile [497] or they are fixed in particular locations [498]. Recent work has also shown that enzymes themselves may form phase-separated clusters with a characteristic equilibrium size [499], similar to the size control exhibited by externally maintained droplets discussed in section 5.3.1. As this field becomes more developed, there will surely be many additional phenomena that we are yet to unveil.
We here presented a continuous theory of chemically active droplets, which negelects a lot of molecular details of the involved biomolecules. For instance, even simple molecular effects, like crowding, can affect active droplet [500]. In the future, it will thus be important to study thermodynamically consistent molecular models of chemically active droplets [501].
@@ 6.1 Life cycle of droplets in biological cells
Droplets need to be in the right place to fulfil their function. To control position, cells can use wetting (section 4) to anchor droplets at particular structures, such as the plasma membrane or chromosome loci, so that they remain in place. If droplets are not anchored, cells can reposition them, e.g., using external gradients (section 2.3.2), hydrodynamic fluxes, or active transport [508]. Even more exciting are droplets that swim themselves, either because they self-organise chemical gradients (section 5.5.2) or they exert active stresses [89, 509, 510].
@@ 6.3 Outlook
Beyond exploring how droplets enable cellular function, understanding the physical principles that biology exploits so effectively has application beyond cell biology. For example, it has been hypothesised that phase separation could be implicated in the origin of life by enabling the formation of basic protocells [148]. Indeed, droplets provide well-defined entities that grow, move, and divide by metabolising material from their environment [523, 524]. On the more theoretical side, chemically active droplets comprise an exciting class of non-equilibrium systems [89], with similarities to proliferating active matter [525] and general dissipative structures [526]. Finally, understanding droplet regulation in biological cells will allow us to leverage similar principles in designing novel synthetic systems [65, 70].
@@ A Compressible multi-component fluid
where we introduced the exchange chemical potentials $\hat{\bar{\mu}}_{j}=\hat{\mu}_{j}-m_{j}m_{0}^{-1}$ relative to solvent and the mobility

### v4

@@ 5.5 More complex chemically active droplets
In this section, we have presented a continuous theory of binary chemically active droplets, which neglects many molecular details of the biomolecules involved. Including such details, even for simple effects such as molecular crowding [482], can lead to additional complex behaviour [483]. Similarly, when many interacting species are involved, we expect to find further interesting effects. In the future, it will thus be important to study thermodynamically consistent models of chemically active multicomponent mixtures, as well as how these mixtures interact with their environment. Since this topic represents a relatively new field of research, we highlight only a few recent discoveries in this direction below.
@@ 5.5.2 Self-organized drift
The combination of chemical turnover and diffusion generically creates concentration gradients, which can move particles and droplets by diffusiophoresis [485, 454, 189, 190, 486]. Concentration gradients also develop around chemically active droplets (figure 5.3), implying that they can influence the positioning of other droplets. For instance, externally maintained droplets naturally repel each other, which stabilises their hexagonal stationary pattern. Their repulsion follows directly from the electrostatic analogy discussed in section 5.3.2.
@@ 5.5.3 Active multi-component mixtures
Active multi-component mixtures can exhibit various instabilities and patterns, ranging from simple Turing patterns [490, 445, 446, 472] to systems with multiple length scales [491], to more complex spatiotemporal behaviour [492, 493, 494]. Particularly interesting is the case where phase separation interferes with chemical reactions [495, 496], so that behaviours beyond those described above can emerge. For instance, droplets may now be stable even in the presence of passive reactions, which is impossible in binary systems [442]. In this case, internally maintained droplets do not fill the entire system, although their size typically still scales with system size [451, 435]. Externally maintained droplets, however, can either scale with system size or exhibit the size control discussed in section 5.3 [458, 497]. In both of these cases, the droplet size distribution relaxes to a monodisperse (and universal) distribution at late times [497]. Externally maintained droplets can also assume shell-like structures [453, 498], essentially because chemical reactions deplete the inside of the droplet so that concentration (figure 5.3A) drops below the spinodal concentration and a bubble emerges. Chemically active droplets may be further controlled by enzyme positioning, either because they are mobile [499] or they are fixed in particular locations [500]. Recent work has also shown that enzymes themselves may form phase-separated clusters with a characteristic equilibrium size [501], similar to the size control exhibited by externally maintained droplets discussed in section 5.3.1. As this field becomes more developed, there will surely be many additional phenomena that we are yet to unveil.
@@ 6.1 Life cycle of droplets in biological cells
Droplets need to be in the right place to fulfil their function. To control position, cells can use wetting (section 4) to anchor droplets at particular structures, such as the plasma membrane or chromosome loci, so that they remain in place. If droplets are not anchored, cells can reposition them, e.g., using external gradients (section 2.3.2), hydrodynamic fluxes (section 3.2.1), or active transport [508]. Even more exciting are droplets that swim themselves, either because they self-organise chemical gradients (section 5.5.2) or they exert active stresses [89, 509, 510].
@@ 6.3 Outlook
Beyond exploring how droplets enable cellular function, understanding the physical principles that biology exploits so effectively has application beyond cell biology. For example, it has been hypothesised that phase separation could be implicated in the origin of life by enabling the formation of basic protocells [148]. Indeed, droplets provide well-defined entities that grow, move, and divide by metabolising material from their environment [523, 524]. On the more theoretical side, chemically active droplets comprise an exciting class of non-equilibrium system [89], with similarities to proliferating active matter [525] and general dissipative structures [526]. Finally, understanding droplet regulation in biological cells will allow us to leverage similar principles in designing novel synthetic systems [65, 70].
@@ A Compressible multi-component fluid
where we have defined exchange chemical potentials $\hat{\bar{\mu}}_{j}=\hat{\mu}_{j}-m_{j}m_{0}^{-1}\hat{\mu}_{0}$ relative to solvent and mobilities